Transition from morning work (*:responsible person)
- ICRR network is still not stable.
- RT model for SRM is down from yesterday. It will be fixed in the morning. If need to look at, it will be afternoon.
Report of work (*:responsible person)
- MIR: First contact was put on MC mirrors yesterday. It will be removed tomorrow.
- IOO: MZ common mode servo was installed.
Work plan (*:responsible person)
- PEM: One of the out put of SR3 DAC was connected to PEM signal.
- VAC: IXC is being vented. Fabian will look the drift and compensate if needed.
- IOO: Noise measurement for optical fiber for green.